OpenStudy (anonymous):

Calculate this integral \[\int\limits_{0}^{1} \frac{ x^2 +1 }{ x^3 +3x+2 } d \theta\]

OpenStudy (mimi_x3):

partial fractions \[ \frac{x^2+1}{ (x+1)(x+2)} = \frac{A}{(x+1)} + \frac{B}{x+2}\]

OpenStudy (mimi_x3):

\[ A(x+2) + B(x+1) = x^2+1

OpenStudy (mimi_x3):

\[ A(x+2) + B(x+1) = x^2+1\]

OpenStudy (mimi_x3):

\[ x = -2\] \[ A(-2+2)+ B(-2+1) = (-2)^2+1)\] \[-B = 5 ->> B = -5\]

OpenStudy (mimi_x3):

\[ x = -1 \] \[ A(-1+2) + B(-1+1) = (-1)^2+1\] \[ A = 2\]

OpenStudy (mimi_x3):

therefore, \[ \frac{2}{(x+1)} - \frac{5}{(x+2)}\]

OpenStudy (mimi_x3):

you can integate it now.

OpenStudy (mimi_x3):

\[ 2\int \frac{1}{(x+2)} dx - 5\int\frac{1}{x+2)}dx\]

OpenStudy (anonymous):

I'm not sure this works, I think maybe you couldn't see that the power that the x on the bottom row is raised to is 3 not 2, so can't factorise it into x+1 and x+2

OpenStudy (mimi_x3):

oh pellet.

OpenStudy (anonymous):

sorry!

OpenStudy (mimi_x3):

long division

OpenStudy (mimi_x3):

nevermind u = x^3+ 3x +2

OpenStudy (mimi_x3):

du / dx = 3x^2 + 3

OpenStudy (mimi_x3):

\[ \int \frac{x^2+1}{x^2+3x+2}dx\] \[ \int \frac{x^2+1}{u}* \frac{du}{3(x^2+1)}\] \[ = \frac{1}{3} \int \frac{1}{u}du\]

OpenStudy (mimi_x3):

\[ = \frac{1}{3} ln(u) +x\]

OpenStudy (mimi_x3):

\[ \frac{1}{3} ln(3x^2+3) + C \]

OpenStudy (mimi_x3):

the x, isnt ment to be there in the second one.

OpenStudy (mimi_x3):

then sub in the limits.